OpenStudy (anonymous):

Help, test coming up! In a frictionless collision, a civic (mass 1200 kg) moving at 23.08 m/s hits resting sentra (mass 1200 kg). The civic moves 30 degrees above the horizontal. What is the speed of the sentra after the collision? I would really appreciate if someone explained this to me!!!!!

OpenStudy (anonymous):

Momentum is conserved in both the x and y directions. x: (1200)(23.08) = 1200 v1 cos(30) + 1200 v2 cos@ y: 0 = 1200 v1 sin(30) + 1200 v2 sin@ Kinetic energy also conserved (1/2) 1200 (23.08)^2 = (1/2) 1200 v1^2 + (1/2) 1200 v2^2 where v1,v2 are civic and sentra velocities after collision, and the sentra goes off at angle 2 to the x-axis. You have three equations and three unknowns [v1, v2, @]. |v2| is the speed of the sentra Hard, but not impossible.

OpenStudy (anonymous):

can we assume the angle is 60 degrees though? because the collision causes them to be perpendicular, and thhe top angle is 30 deegrees

OpenStudy (anonymous):

also how do you know kinetic energy is conserved?

OpenStudy (anonymous):

"frictionless" hints no energy losses, and the cars do not end up entwined, so it seems they do want to assume energy is conserved. Without that, you have no third equation. Does the problem say the are perpendicular after collision? What if it were head-on or only grazing incidence? I would not assume perpendicularity.

OpenStudy (anonymous):

forgive my ignorance, by thought there was only no energy loss if it directly said it was elastic?

OpenStudy (anonymous):

Ideally, they would tell you "elastic." Without either "elastic" or "perfectly inelastic," you are up in the air. I think "frictionless" is being used as "elastic." But who knows?